One call option with a maturity of 3 months and an exercise price of

15,000 is purchased at 4,000 Won and two call options with an exercise

price of 17,500 at the same maturity are sold for 2,000 Won and a call

option with the same maturity and the strike of 20,000 is longed at 500

Won. What is the profifit or loss of this option investor if the underlying

asset price reaches 18,000 Won at maturity?
Answer #1

Sol:

Maturity price = 18,000 Won

Purchase price of one 15,000 Call option = 4,000 Won

Sold price of two 17,500 Call option = 2,000 Won x 2 = 4000 Won

Purchase price of one 20,000 Call option = 500 Won

To determine profit or loss to the investor, if the underlying asset price reaches 18,000 Won at maturity as follows:

1) Purchase price of one 15,000 Call option = 4,000 Won.

Breakeven price = 15,000 + 4,000 = 19,000 Won

Profit/Loss = Maturity price - Breakeven price

Loss = 18,000 - 19,000 = -1,000 Won

2) Sold price of two 17,500 Call option = 2,000 Won x 2 = 4,000 Won

Profit/Loss = 4000 - (Maturity price - option price)

Profit/Loss = 4000 - (18,000 - 17,500)

Profit = 4000 - 500 = 3500 Won

3) Purchase price of one 20,000 Call option = 500 Won

Loss = Premium paid for out of the money call as maturity price is 18,000 = -500 Won

Profit on the option investment will be = 3500 - 1000 - 500 = 2000 Won

Therefore profit to the investor, if the underlying asset price reaches 18,000 Won at maturity will be 2000 Won.